OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 4x – 6y = 30 3x – 5y = 24 (Points : 20)

OpenStudy (anonymous):

4x – 6y = 30 ---eqn(I) 3x – 5y = 24 ---eqn(II) multiply eqn(I) with 3 and eqn(II) with 4 3(4x – 6y = 30) 12x-18y=90 ----eqn(III) 4(3x – 5y = 24) 12x-20y=96 -----eqn(IV) subtract eqn(IV) from eqn(III) 12x-18y=90 -(12x-20y=96) --------------- 2y= -6 so y=-3 put this value in any of the given eqns and get value of x 4x – 6y = 30 4x-6(-3)=30 4x+18=30 3x=12 x=3 so the solution set is (3,-3)